Startup Night 8 - 20140906 UT

Observers: JHill, BRothberg (LBTO) DMiller (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: JCGuerra
Telescope Support: JUrban (LBTO)
Software Support: MDeLPena, KSummers
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby


Clouds cleared off enough to open by 3 UT.

Got some final wake-up observations with LBC and identified a bug in the code. Clouds got too thick for LBC, so switch to active optics calibrations with AGW1 at 6:30. Made good progress on crosstalk with AGW1 and zernike calibrations with MODS1 once clouds thinned out. Switched to MODS1 observations at 10:10 and got a standard. Closed by humidity an hour later.


James and Conrad fixed both LBC-Red Shutter and the LUCI1 communications issues.

Michele and Dan gave us a new version of ECS and ECSGUI.


Had a visit by Eagle Gamma -- -- and showed him the telescope around sunset. It took him 6 hours to bicycle up from the basecamp. Humidity has been sitting at 92% for a while.

02:55 Humidity is dropping so we will try to open.


03:03 Open to a few clouds and rapidly dropping humidity. 73% T=10.7, D=6.0, wind light out of SE. Chamber is close to ambient, but mirrors are still 3-4 degC warm.

OSU monitor N6503

03:05 OSU_monitor_N6503focus.ob


03:17 OSUmonitorN6503.ob -- telescope is drifting so that guiding is struggling to keep up -- why? The rotator speed is wrong at EL=51. Stopped OB after the first image.

03:25 Switch PCS back to the old pointing kernel PCS.2014C

03:28 OSUmonitorN6503.ob again

Call Michele

Call Kellee

Could it be a hidden rotator error not reported on UI?

03:39 Turn off / on LBC software crashed twice during turn on

03:40 Back to PCS.tpk2.3 the new pointing kernel

03:46 OSU_monitorN6946.ob without focussing - problem persists

Kellee sees in the log that LBC thinks it doesn't have a rotator trajectory.

03:51 lbckill/lbcstart to get rotator diagnostics (flag in config file)

03:53 OSU_monitorN6946.ob without focussing

03:59 lbckill/lbcstart to get version of 2-Sep-2014 lbc software

04:02 OSU_monitorN6946.ob without focussing -- rotator is OK

04:08 OSU_monitor_N6503focus.ob

04:09 DOFPIA blue struggling with focus convergence

04:17 OSUmonitorN6503.ob -- and blue guide star looks not well focussed. SX Primary Actuator 423 is disabled. Image has a bit of trefoil. Red images are beautiful.

Arp314 Dithered Observation

04:33 Arp314focusdual

04:36 DOFPIA (stopped manually at EL=38)

04:42 Arp314dual for dithered deep observation -- (as deep as you can with a gibbous moon) -- seeing is rather puffy down here at EL=38. shutter closed on red from OB.

04:46 replayed editted Arp314dua

04:49 DOFPIA

04:51 Arp314dual_back.ob was OK until some clouds near the moon drifted across us about halfway through the OB.

05:21 Arp314_dual_rotate_VI PA=90 - still fighting thin clouds -- I guess we'll call them sky flats with 55 counts in Blue.

NS 2060 Chiron

05:32 Hijacked LBC preset to Chiron - rate ~8 arcsec per hour - loads of moonlit clouds, but these are short exposures.

05:38 WT10_323 in a part of the sky where there is less cloud.

05:39 SX Primary Panics during preset - why?

05:55 WT10_323

Clear Active Optics

06:00 DOFPIA, /FIRST -- too much cloud

DOFPIA needs to wait slightly longer for the next OB when no pupils are found. But, shouldn't DOFPIA have exited when no pupils found? The exit only works on the first iteration. I added 2 sec of WAIT to before it plays the OB.

Reconstructor Tests

06:17 Reconfigure to LUCI1 as too many clouds for LBC. Doug is on the Polycom from home 3-wayed to N525.

06:39 LUCI1 on left for AGW1 and LBC-Red on right.


06:43 ACTIVE preset to BS9188 with Doug's new reconstructor -- failed co-pointing limit even with pseudomono mode Can't do pseudo mono when authorized as ARGOS -- targets are too far apart?

06:46 preset to BS9188 with Doug's new reconstructor
2014-09-06T06:57:49.916793+00:00 tcs4 LBT_left_GCS: WFS: triggerOldWFS: Subimage could not be created, viewing full WFS frame in GUI instead 
2014-09-06T06:57:49.916799+00:00 tcs4 LBT_left_GCS: WFS: triggerOldWFS: createSubImage(386, 208, 254, 254) failed.
2014-09-06T06:57:49.933888+00:00 tcs4 LBT_left_GCS: WFS: execute: Source Extractor will be used.
2014-09-06T06:57:49.933899+00:00 tcs4 LBT_left_GCS: WFS: execute: Dynamic reconstruction matrix will be used.
2014-09-06T06:57:49.933905+00:00 tcs4 LBT_left_GCS: WFS: execute: supposed to attempt new WFS approach but Sextractor object not instantiated!
2014-09-06T06:57:49.933909+00:00 tcs4 LBT_left_GCS: WFS: execute: using old style WFS instead ...
2014-09-06T06:57:49.933914+00:00 tcs4 LBT_left_GCS: WFS: triggerOldWFS start.

Change back to old reconstructor

07:00 preset to BS9188 with
2014-09-06T07:01:49.751228+00:00 tcs4 LBT_left_GCS: WFS: execute: thread started.
2014-09-06T07:01:49.760304+00:00 tcs4 LBT_left_GCS: WFS: execute: Source Extractor will be used.
2014-09-06T07:01:49.760344+00:00 tcs4 LBT_left_GCS: WFS: execute: Dynamic reconstruction matrix will be used.
2014-09-06T07:01:49.760370+00:00 tcs4 LBT_left_GCS: WFS: execute: supposed to attempt new WFS approach but Sextractor object not instantiated!
2014-09-06T07:01:49.760397+00:00 tcs4 LBT_left_GCS: WFS: execute: using old style WFS instead ...
2014-09-06T07:01:49.760431+00:00 tcs4 LBT_left_GCS: WFS: triggerOldWFS start.

07:03 Stop and restart GCSL

07:04 preset to BS9188 with old reconstructor -- worked OK

StopWFSing to install new reconstructor.

07:07 preset to BS9188 with Doug's new reconstructor

07:10 Clear Active Optics M1 after Z6 runaway.

07:20 ACTIVE preset to WT10_351 gs=0 to get something brighter.

07:28 Back to old reconstructor for crosstalk offset test -- clouds part and WT10_351 is too bright.

07:33 setHotspot left -x 261 -y 211 (was 260,214)

07:34 setHotspot left -x 259 -y 211 THIS IS THE NEW HOTSPOT -- added to LUCI_L.cfg


Doug running test of zernike_collect with cross.list

offset 25 0 DETXY ABS wrong way

07:46 offset -25 0

07:54 back to BS9188 but still to much cloud

LFBG crosstalk

data in /home/lbto/data/20140906/lfbg_cross/

07:59 AO56, offset -25 0


zernike_collect, exp=6 with cross.list and default limit=300
rw-rw-r--. 1 root      tcs          4303 Aug 26 16:42 psf.conf
-rw-r--r--. 1 cbiddick  domain users 4303 Jun 18 23:31 psf.conf.20140618
-rw-r--r--. 1 telescope domain users 4318 Sep  6 08:19 psf.conf.cross
-rw-r--r--. 1 telescope domain users 4303 Jun 14 06:05 psf.conf.orig
[telescope@obs2 etc]$ cp -p psf.conf psf.conf.20140826
[telescope@obs2 etc]$ cp -p psf.conf.cross psf.conf
cp: overwrite `psf.conf'? y
08:22 stopped the first zernike_collect which was with old crosstalk terms (Doug will move above data to sub-directory.)

08:23 stop and restart PSFL to read the crosstalked version of psf.conf

08:25 zernike_collect, exp=12, limit=400 with cross.list   seeing improved to 0.6 on guider and fewer clouds.

---++++ MODS1 Zernikes

08:50 switching over to MODS1 and LBC-Red
cp -p psf.conf.20140826 psf.conf 
cp: overwrite `psf.conf'? y
cp: preserving times for `psf.conf': Operation not permitted

to remove crosstalk from psf.conf
08:51 stop and restart PSFL 09:02 ACQUIRE preset
09:04 ACTIVE preset to AO56, but its too bright when there are no clouds
09:09 ACTIVE preset to BS9107
09:13 offset -20 0 09:15 offset -25 0
09:18 offset -30 5 09:20 offset 30 5 (is it really rotated 90 compared to AGW1?)
09:23 offset 5 30 (probe failed to moved)
09:25 offset 5 -30 stopWFSing data in /home/lbto/data/20140906/ldg_zern/ 09:27 zernike_collect, limit=400 with full.list with default 15 sec exposure.
10:10 zern_collect completed - humidity up to 88% again

Feige 110

UT 10:15 - acquiring Feige 100 using 0.6" slit (mods1r.20140906.003)

UT 10:18 - modsDisp did not show the thru-slit image (mods1r.20140906.003) but it was saved and appeared using modsAlign

UT 10:22 - feige 110 slightly offset, sending +0.123" detx preset, sending -0.246" detx

UT 10:26 - feige 110 - 0.6" slit (mods1r.20140906.0007/mods1b.20140906.0003) seeing 1.2" guider -nonphotometric conditions

UT 10:28 - red IMCS lock fail - retrying - worked - airmass 1.753

FF0139+0115 (rothberg)

UT 10:38 - moving to target - when moving here from feige110 odd "thump" sound occurred in enclosure

UT 10:42 - acquiring (mods1r.20140906.0010)

UT 10:51 - aligning in longslit

UT 10:53 - aborted thru slit image - humidity hit 95% frown, sad smile


UT 10:56 - Closing Enclosure

UT 11:05 - MODS1 put to sleep
Topic revision: r4 - 06 Sep 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback